Circular No. 6002

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA OPHIUCHI 1994 R. M. Wagner, Ohio State University; K. M. Vandlandingham, Ari-

zona State University; and N. King, New Mexico State University, report that optical spectra (range 367–723 nm, resolution 0.36 nm) were obtained on June 4.31 UT with the 4.5-m Multiple Mirror Telescope at the F. L. Whipple Observatory. The spectra exhibit broad permitted emission lines due to the Balmer series of hydrogen and Fe II superposed on a flat continuum. The FWHM of Hα and Hβ emission is ∼ 2500 km/s and the FWZI is ∼ 5800 km/s. The line profiles consist of at least four distinct components. The spectra confirm the identification as a fast Fe II-class nova in its early decline stage.

S. Benetti, European Southern Observatory (ESO), reports that inspection of a preliminarily-reduced CCD spectrum (range 370–700 nm, resolution 1.2 nm) obtained on June 4.24 UT with ESO 3.6-m telescope (+ EFOSC1) confirms that the object is a nova. The spectrum is dominated by intense emission lines of hydrogen and Fe II. The FWHM of Hα, Hβ, and Hγ, corrected for the instrumental resolution, is ∼ 2500 km/s.

S. D. Van Dyk, University of California at Berkeley, obtained a spectrum (range 470–720 nm, resolution 0.4 nm) of N Oph 1994 on June 4 UT with the Kitt Peak 2.1-m reflector (+ Goldcam spectrograph). G. S. Stringfellow (Penn State University) and A. V. Filippenko (University of California at Berkeley) have examined the uncalibrated spectrum; they report that broad emission lines are present, by far the strongest being Hα. Prominent [O III] (495.9, 500.7 nm) emission is evident, along with many weaker forbidden lines such as [O I] at 630.0 and 636.4 nm. He I at 587.6 nm is narrower and exhibits a P-Cyg profile. Structure is visible in most of the emission lines, suggesting that the ejecta are clumpy. A spectrum centered on Hα, obtained by G. E. Albright (University of Virginia) at the coudé feed telescope, confirms the presence of at least four distinct components.

A. C. Gilmore reports the following preliminary photometry from Mt. John (reference star from Cousins E746; uncertainty ∼ 0.02 in each color except U − B, where the error may be 0.05): June 4.585 UT, V = 8.49, U − B = −0.25, B − V = +0.52, V − R_{c} = +0.66, V − I_{c} = +1.29.

Magnitude estimates from T-Max film from K. Kanatsu, Matsue, Shimane, Japan: May 30.585 UT, [9.5; June 2.653, 7.5.

1994 June 4 Daniel W. E. Green